(a) Boise Company manufactures and sells three products: Good, Better, and Best. Annual fixed costs are $3,315,000, and data about the three products follow. Good 30% $250 Better 50% $350 Best 20% $500 Sales mix in units Selling price Variable cost 100 150 250
1. Determine the weighted-average unit contribution margin 2 Determine the break-even volume in units for each product 3. Determine the number of units that must be sold for each product to obtain a profit for the company of $234,000.
(a) Adelaide Pty Ltd manufactures and sells three products: Xeno, Yanko and Zengo. Annual fixed costs are $1,360,000. Xeno Yanko Zengo Sales mix in units 40% 35% 25% Selling price Variable costs $200 $600 $800 $100 $280 $320 Required: (i) Calculate the weighted average unit contribution margin, assuming a constant sales mix
ii)Determine the break even volume in units and in sales dollars for each product? Assume a constant sales mix ii) Determine the number of units of each product that must be sold to obtain a profit; of $544,000 before tax? Assume a constant sales mix.
Answer #1

Boise Company

1.

Good Better Best
Sales mix (A) 30% 50% 20%
Selling price per unit (p.u.) (B) 250 350 500
Variable cost p.u. (C) 100 150 250
Contribution margin p.u. (D = B - C) 150 200 250
Weights assigned to each product (A * D) 45 100 50

Weighted average contribution margin p.u. = 45+100+50
$ 195

2. Break even sales in units = total fixed cost / average contribution margin
= 3315000/195
= 17,000 units

Break even point for each product:

Good = 17,000 * 30% = 5,100 units

Better = 17,000 * 50% = 8,500 units

Best = 17,000 * 20% = 3,400 units

(to get break even unit in $, multiply the number of units with the respective selling prices)


3. Required number of units to be sold = (total fixed cost +desired profit) / average contribution
= (3315000+234000) / 195
= 18,200 units

Number of units for each product:

Good = 18,200*30% = 5,460 units

Better = 18,200*50% = 9,100 units

Best = 18,200*20% = 3,640 units

The next question is the same as the first question, but with different figures. So, please solve it using the same process as explained above.